The Cost of Equity and Flotation Costs

Messman Manufacturing will issue common stock to the public for $35. The expected dividend and the growth in dividends are $2.75 per share and 6%, respectively. If the flotation cost is 9% of the issue's gross proceeds, what is the cost of external equity, re? Round your answer to two decimal places.
%
Answer #1

Solution:

Price of stock = $35

Flotation cost = 9%

Net proceeds = $35*(1-9%) = $31.85

Next expected dividend = $2.75

Growth rate = 6%

Cost of external equity ( re) =( Next expected dividend / Net proceeds ) + growth rate

Cost of external equity ( re) = (2.75/31.85)+6%= 14.63%